OpenStudy (anonymous):

Check my answers? I'm taking a little practice quiz and I want to be sure I've got my stuff down pat before moving on to the real thing.

OpenStudy (anonymous):

\[\frac{ z ^{3} y ^{4} }{ z ^{5}y }\]

OpenStudy (anonymous):

and my answer is z^2 y^3

OpenStudy (mathstudent55):

The y^3 is correct, but not the z part.

OpenStudy (anonymous):

is the z part supposed to be negative then?

OpenStudy (amistre64):

depends on the instructions

OpenStudy (amistre64):

if negative expos are fine, then sure; otherwise we need to place it properly above/below the fraction bar

OpenStudy (mathstudent55):

If you can leave it as a fraction, then both exponents will be positive, but some of it will be in the denominator.

OpenStudy (anonymous):

It just says simplify so I suppose I should simplify the negative exponent

OpenStudy (mathstudent55):

I'd leave it as a fraction with positive exponents.

OpenStudy (anonymous):

so then 1/z^2 y^3

OpenStudy (mathstudent55):

You know the rule: \(\large \dfrac{a^m}{a^n} = a ^{m - n} \) ? Apply it to the z's and to the y's. Then change the one with a negative exponent into a fraction.

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

can you check my next one?

OpenStudy (mathstudent55):

\(\large \dfrac{ z ^{3} y ^{4} }{ z ^{5}y } = z^{-2}y^3 = \dfrac{1}{z^2} \times y^3 = \dfrac{y^3}{z^2} \)

OpenStudy (mathstudent55):

Sure.

OpenStudy (anonymous):

oh wait so it's not 1/z^2 y^3 it's y^3/z^2?

OpenStudy (mathstudent55):

They both mean the same, but \(\dfrac{y^3}{z^2}\) is simpler.

OpenStudy (anonymous):

so this one is write (3a)^2 without exponents...and ummm the only thing I can come up with is 9a...I don't remember learning this... and then the second part of that is Fill in the blanks (3a)^2 = blank a ^blank...which I don't think I understand what they are asking me to do

OpenStudy (mathstudent55):

Ok. I'll explain the rule to you.

OpenStudy (mathstudent55):

Here is the rule as a statement in English: When you raise a product to a power, raise every factor of the product to the power.

OpenStudy (mathstudent55):

Here is the rule as a mathematical expression: \(\large (ab)^n = a^nb^n\)

OpenStudy (mathstudent55):

Notice the mathematical expression just above. It is very similar to your problem. If you raise the product ab to power n, you must raise a to the n power and multiply it by b to the n power. In your case, the product is 3a, and the power is 2.

OpenStudy (anonymous):

ok...so let me figure this out...so I have to raise 3^2 and a^2? which gives me me...9a^4because I add the exponent on the a to the one outside the parentheses? (if I sound stupid forgive me I am just trying to figure this out)

OpenStudy (mathstudent55):

No. You started out correctly. \(\large (3a)^2 = 3^2a^2\) Now you can actually calculate \(3^2\) because that is simply \(3 \times 3\). That means now you have \(9a^2\)

OpenStudy (mathstudent55):

You only add exponents when you multiply two powers with the same base. For example, \(4^5 \times 4^7 = 4^{5 + 7} = 4^{12}\)

OpenStudy (anonymous):

ahhh so it stays at 9a^2 ok that makes sense...so what are they talking about this fill in the blank thing in the second part? I still don't get that...

OpenStudy (anonymous):

oh wait...no that part IS the fill in the blank part...the second part is the one where I have to rewrite it without exponents.

OpenStudy (mathstudent55):

The second part is "blank a ^blank", right?

OpenStudy (mathstudent55):

The second part is simply: \(\large (3a)^2 = 9a^2\)

OpenStudy (anonymous):

right the first part is the part where I have to write it without exponents.

OpenStudy (mathstudent55):

For the first part, since you must write it without exponents, you must write the meaning of a^2 as a product. What does a^2 mean as a product without using exponents?

OpenStudy (anonymous):

a times a

OpenStudy (mathstudent55):

Exactly. That means, \((3a)^2 = 3^2 a^2 = 9aa = 9 \times a \times a \)

OpenStudy (anonymous):

so they just want me to give the equation...geez my class is weird

OpenStudy (mathstudent55):

Maybe their idea is to make sure you understand what exponents mean.

OpenStudy (anonymous):

(2z^2)^4 simplify without parentheses and my answer was 16z^8

OpenStudy (mathstudent55):

Excellent.

OpenStudy (anonymous):

really?!?!?

OpenStudy (anonymous):

I got it?

OpenStudy (mathstudent55):

I wouldn't say "Excellent, you're wrong." That would be mean. I wrote Excellent to mean you did excellent work and are correct.

OpenStudy (mathstudent55):

Yes, you got it. I don't know why you're so surprised. You sound pretty smart to me. You understood quickly the rule of exponents I showed you and applied it yourself correctly.

OpenStudy (anonymous):

Sorry, I felt stunned...I'm usually so bad at getting this stuff...and my teacher this year has been very hard on me

OpenStudy (mathstudent55):

Keep studying and asking questions. That's how you learn. I think you're on the right track.

OpenStudy (anonymous):

(a^3/-3b)^2 write without parentheses a^6/9b^2

OpenStudy (mathstudent55):

Correct & excellent, and don't be stunned.

OpenStudy (anonymous):

I really have to thank you for helping me get my confidence :)

OpenStudy (mathstudent55):

You are welcome. Keep practicing and asking questions.

OpenStudy (anonymous):

Ok I think I've almost got this one but I got a little stuck... find two consecutive whole numbers that sqrt65 lies in between and ummm so far I've got that 65 falls in the interval between 64 and 81 and it falls between 8 and 9 and then when I get the difference of 65 and 64 I get 1 and then I get the difference of 81 and 65 and I get 16...which is not 1/16 correct?...but now I'm stuck...

OpenStudy (mathstudent55):

Yes, \(\sqrt{65} \) does lie between 8 and 9.

OpenStudy (mathstudent55):